# Tilting representations of finite groups of Lie type

### v1

@@ 4.2. A conjecture of Dudas–Malle
The goal of this section is to provide an answer to a conjecture of Dudas and Malle [DM14, Conjecture 1.2]. The following hypotheses are in force in this section
  1. $(i)$.
    All the Deligne-Lusztig varieties of $(\mathbf{G},\mathrm{F})$ are affine.
  2. $(ii)$.
    The Kazhdan-Lusztig basis and the $\ell$-Kazhdan-Lusztig basis coincide. By Lemma 2.6.3 this is guaranteed whenever $\ell$ is large enough (depending only on $\mathbf{G}$ and not on the Frobenius).
@@
We denote by $\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{Weil}}$ the category of pairs $(\rho,\alpha)$ where $\rho\in\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})$ and $\alpha$ is an isomorphism $\mathrm{F}^{\delta,*}\rho\simeq\rho$. For $\lambda\in\overline{\mathbb{Q}}_{\ell}$ and $(\rho,\alpha)\in\mathrm{D}_{c}(\mathrm{pt}/\mathbf{G}^{\mathrm{F}},\overline{\mathbb{Q}}_{\ell})^{\mathrm{Weil}}$, we denote by
$\rho[\lambda]$
$\rho[\overline{\lambda}]$
the direct sum
$\oplus_{\lambda}\rho[\lambda]$
indexed by all $\lambda\in\overline{\mathbb{Z}}_{\ell}$ which reduce to $\overline{\lambda}\in\overline{\mathbb{F}}_{\ell}$.
@@
$d(\mathrm{ch}(\mathrm{IC}_{w,1}^{\mathrm{mix}}[\lambda]))$
@@
$T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=T_{w,1}^{\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{\chi_{\ell}\neq 1}T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{v<w,\chi_{\ell}}T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\oplus_{n_{v,w,\chi_{\ell}}}}$
@@
$T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}\oplus\bigoplus_{\chi_{\ell}}T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}\oplus\bigoplus_{v<w,\chi_{\ell}}T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix},\oplus_{n_{v,w,\chi_{\ell}}}}.$
@@
$\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])[\overline{\lambda}]=\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]\oplus\bigoplus_{\chi_{\ell}}\mathrm{ch}(T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]\oplus\bigoplus_{v<w,\chi_{\ell}}\mathrm{ch}(T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix},\oplus_{n_{v,w,\chi_{\ell}}}})[\overline{\lambda}].$
Note that the unipotent part of this representation is $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]$. Since mixed tilting sheaves are filtered, as sheaves with Weil structure, by mixed standard sheaves, it follows that the only eigenvalues of Frobenius appearing in $\mathrm{ch}(T_{w,\chi}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})$ are eigenvalues of Frobenius appearing on the cohomology of Deligne-Lusztig varieties. Those eigenvalues are known to be of the form $q^{\frac{i}{2}}\zeta$ where $\zeta$ is a root of $1$ and $i\in\mathbb{Z}$ by [Lus84, Theorem 3.8.1] . It follows that the Weil structure on $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])$, which is an isomorphism
@@
It is then enough to show that when $\lambda=q^{\frac{i}{2}}\zeta$, we have
$d(\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}[\lambda]))=\mathrm{ch}(\mathrm{IC}_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\lambda].$
It follows from Corollary 4.1.4 that we have
$d(\sum_{\zeta}\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}[q^{\frac{i}{2}}\zeta]))=\sum_{\zeta}\mathrm{ch}(\mathrm{IC}_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[q^{\frac{i}{2}}\zeta].$
From [Lus84, Theorem 3.8.1], attached to every irreducible unipotent representation $\rho$, there is a root of $1,\zeta_{\rho}$ such that for any Deligne-Lusztig variety $X(w)$, the eigenvalues of Frobenius on the $\rho$-isotypic component of the cohomology of $X(w)$ are all of the form $\zeta_{\rho}q^{N}$ for various half integers $N$. This property holds for any object of the form $\mathrm{ch}(A)$, where $A\in\mathcal{O}^{\mathrm{DL},\mathrm{unip},\mathrm{mix}}$ as the standard sheaves (whose images are the cohomology complexes of the Deligne-Lusztig varieties) generate this category.
This root of unity satisfies
$\zeta_{\pm d(\rho)}=\zeta_{\rho},$
where ±$\pm d(\rho)$ is the Alvis-Curtis dual of $\rho$, i.e. this root of $1$ is invariant under the duality. By [Lus78, 3.33] (see also [GM03, Lemma 4.2]), it follows that this root of $1$ is constant on Harish-Chandra series, which implies our claim. It follows from these properties that we can isolate each summand corresponding to a given $\zeta$ in the previous sums, in particular we get
$d(\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}[q^{\frac{i}{2}}\zeta]))=\mathrm{ch}(\mathrm{IC}_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[q^{\frac{i}{2}}\zeta].$

### v2

@@ 4.3. A conjecture of Dudas–Malle
The goal of this section is to provide an answer to a conjecture of Dudas and Malle [DM14, Conjecture 1.2]. The following hypothesis is in force in this section :
The Kazhdan–Lusztig basis and the $\ell$-Kazhdan–Lusztig basis coincide.
By Lemma 2.6.3 this is guaranteed whenever $\ell$ is large enough (depending only on $\mathbf{G}$ and not on the Frobenius).
@@
$d(\mathrm{ch}(\mathrm{IC}_{w,1}^{\mathrm{mix}}[\overline{\lambda}]))$
@@
$T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=T_{w,1}^{\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{\chi_{\ell}\neq 1}T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell}}\oplus\bigoplus_{v<w,\chi_{\ell}\neq 1}T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\oplus{n_{v,w,\chi_{\ell}}}}$
@@
$T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}]=T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}\oplus\bigoplus_{\chi_{\ell}\neq 1}T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}}\oplus\bigoplus_{v<w,\chi_{\ell}\neq 1}T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix},\oplus{n_{v,w,\chi_{\ell}}}}.$
@@
$\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])[\overline{\lambda}]=\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]\oplus\bigoplus_{\chi_{\ell}\neq 1}\mathrm{ch}(T_{w,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]\oplus\bigoplus_{v<w,\chi_{\ell}\neq 1}\mathrm{ch}(T_{v,\chi_{\ell}}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix},\oplus{n_{v,w,\chi_{\ell}}}})[\overline{\lambda}].$
The unipotent part of this representation is $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})[\overline{\lambda}]$. Since mixed tilting sheaves are filtered, as sheaves with Weil structure, by mixed standard sheaves, it follows that the only eigenvalues of Frobenius appearing in $\mathrm{ch}(T_{w,\chi}^{\overline{\mathbb{Q}}_{\ell},\mathrm{mix}})$ are eigenvalues of Frobenius appearing on the cohomology of Deligne–Lusztig varieties. By Theorem 4.2.1, these eigenvalues of Frobenius are of the form $q^{\frac{i}{2}}\zeta$ where $i\in\mathbb{Z}$ and $\zeta\in\overline{\mathbb{Q}}_{\ell}^{\times}$ are roots of unity. It follows that the Weil structure on $\mathrm{ch}(T_{w,1}^{\overline{\mathbb{Z}}_{\ell}}[\frac{1}{\ell}])$, which is an isomorphism
@@
This is exactly Corollary 4.2.4. ∎